Non-Exact Differential Equations. Find the complete solution of ydx + (x + y ^ 2) dy = 0

1
Expert's answer
2022-03-04T05:53:41-0500
"ydx+(x+y^2)dy=0"

"\dfrac{dx}{dy}+\dfrac{1}{y}x=-y"

"I.F.=e^{\int(1/y)dy}=y"

"ydx+xdy=-y^2dy"

"d(xy)=-y^2dy"

Integrate


"\int d(xy)=-\int y^2dy"

"xy=-\dfrac{y^3}{3}+C"

"xy+\dfrac{y^3}{3}=C"